If you mix 140 mL of a 60% acid solution with 200 mL of a 30% solution you can get a 350 mL of a _____% solution.
Help!

Found 2 solutions by checkley71, joyofmath:
Answer by checkley71(8403)

140*.6+200*.3=350x
84+60=350x
140=350x
x=140/350
x=.40 or 40% mix.

Answer by joyofmath(189)

I think there's a typo in your question. If you combine 140 mL of something with 200 mL of something you get 340 mL, not 350 mL.
If you have 140 mL of a 60% acid solution then you've got 60% * 140 mL = 84 mL of acid in the first solution.
If you have 200 mL of a 30% solution then you've got 30% * 200 mL = 60 mL of acid in the second solution,
When you combine the two solutions you have 84 mL + 60 mL = 144 mL of acid and 140 mL + 200 mL = 340 mL of total liquid.
To get the concentration you divide 144 by 340 and get approximately 42.4%